OpenStudy (jpsmarinho):

how to integrate this: 1/(u^2 + 3) du??

OpenStudy (lgbasallote):

are you familiar with trigonometric substitution?

OpenStudy (jpsmarinho):

noo

OpenStudy (jpsmarinho):

but, i need to learn. If you not incomodate with aux me... ^^

OpenStudy (anonymous):

\[\int \frac{du}{u^2+3}=\int\frac{du}{u^2+(\sqrt{3})^2}\]

OpenStudy (jpsmarinho):

Thanks!!